# Inference for Regression with

### v2

@@ A.2 Proofs for Section 3
$\Bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ -\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\Bigg{\|}\to_{p}0.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\left[\begin{array}[]{cc}\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right)\\ \left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}\end{array}\right].$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ +\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right].$ (19)
@@
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha},$
@@
$\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\right)\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}\,.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}Y_{i}\to_{p}\mathbb{E}[\mathbf{q}_{i}Y_{i}]=\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\bm{\psi}.$ (21)
The first result follows from (19) and (21). Note that the matrix on the right-hand side of (19) is bounded below (in Loewner order) by $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ so its inverse is well defined by Assumption 1(iv). The second result follows because $(\mathbf{A}+\bm{\Delta})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\bm{\Delta}\mathbf{A}^{-1}+O(\|\bm{\Delta}\|^{2})$ for $\mathbf{A}$ invertible and $\bm{\Delta}$ small. ∎
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right).$
@@
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii), where
$T_{1,n}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
@@
Assumption 2(i)-(ii) implies $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. As $\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 1$, it follows that $T_{1,n}\to_{p}\mathbf{0}$. For term $T_{2,n}$, note by Lemma 1 that
@@
$=\mathbb{E}\left[\frac{1}{C_{i}}\right]\left(\mathrm{diag}\left(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right]\right)-\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}\right).$ (22)
@@
We conclude that $T_{2,n}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right)$ by (7), (22), and Assumption 2(i)-(ii)
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}\mathbf{0}.$
@@
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii)-(iv), where
$T_{1,n}$ $=\left(\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
@@
Assumption 2(i)-(ii) implies that $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. Moreover, $\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|$, which is bounded in probability by Assumption 2(iv). It follows that $T_{1,n}\to_{p}\mathbf{0}$. For $T_{2,n}$, note that $\mathbb{E}[\mathbf{q}_{i}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}]=\mathbf{0}$ by independence of $\mathbf{q}_{i}$ and $\mathbf{x}_{i}$ conditional on $(C_{i},\bm{\theta}_{i})$. Let $\mathbf{X}_{i}=\mathbf{q}_{i}\left(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right)^{T}$ and $D$ the dimension of $\mathbf{q}_{i}$. Then
@@
First consider the denominator term. By Lemma 5 and condition (7), we have $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}]$. Hence, by Assumption 2(iv),
$\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}]^{-1}.$ (23)
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;(Y_{i}-\hat{\bm{\xi}}_{i}^{T}\bm{\psi})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}(\bm{\xi}_{i}-\hat{\bm{\xi}}_{i})^{T}\bm{\psi}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\varepsilon_{i}=:T_{1,n}+T_{2,n}.$
@@
$T_{1,n}\to_{p}\left[\begin{array}[]{c}-\kappa\,\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$ (24)
@@
$=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|C_{i},\bm{\theta}_{i}\right]\frac{1}{C_{i}}\mathrm{tr}\left\{\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right\}\right]$
$=\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\varepsilon_{i}^{2}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right]\to 0.$
@@
Note that $\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ is finite by Assumption 2(iv)-(v). It follows by the central limit theorem that
$T_{2,n}\to_{d}N\left(\bm{0},\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\right).$ (25)
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right].$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\left(\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-{\bm{\xi}}_{i}\!\!\!\;{\bm{\xi}}_{i}^{T}\right)\\ +\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2})\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=:T_{3,n}+T_{4,n}+T_{5,n}.$
Evidently, $T_{3,n}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ by the LLN. Now consider $T_{4,n}$. By construction, we have
$T_{4,n}=\left[\begin{array}[]{cc}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T})&\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})\mathbf{q}_{i}^{T}\\ \frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}&\mathbf{0}\end{array}\right].$

### v3

@@ A.2 Proofs for Section 3
$\Bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ -\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\Bigg{\|}\to_{p}0.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\left[\begin{array}[]{cc}\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right)\\ \left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\end{array}\right].$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ +\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right].$ (20)
@@
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha},$
@@
$\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\right)\to_{p}\bm{S}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\bm{S}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}\,.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}Y_{i}\to_{p}\mathbb{E}[\mathbf{q}_{i}Y_{i}]=\mathbb{E}\left[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbb{E}\left[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\bm{\psi}.$ (22)
The first result follows from (20) and (22). Note that the matrix on the right-hand side of (20) is bounded below (in Loewner order) by $\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ so its inverse is well defined by Assumption 1(iv). The second result follows because $(\mathbf{A}+\bm{\Delta})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\bm{\Delta}\mathbf{A}^{-1}+O(\|\bm{\Delta}\|^{2})$ for $\mathbf{A}$ invertible and $\bm{\Delta}$ small. ∎
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}]\right).$
@@
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii), where
$T_{1,n}$ $=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
@@
Assumption 2(i)-(ii) implies $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. As $\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 1$, it follows that $T_{1,n}\to_{p}\mathbf{0}$. For term $T_{2,n}$, note by Lemma 1 that
@@
$=\mathbb{E}\left[\frac{1}{C_{i}}\right]\left(\mathrm{diag}\left(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right]\right)-\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}\right).$ (23)
@@
We conclude that $T_{2,n}\to_{p}-\kappa\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}]\right)$ by (7), (23), and Assumption 2(i)-(ii)
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}\to_{p}\mathbf{0}.$
@@
First note that $\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i})^{T}-T_{1,n}-T_{2,n}\right\|\to_{p}0$ by Assumption 2(iii)-(iv), where
$T_{1,n}$ $=\left(\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T}\right)\sqrt{n}\left(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right)\right)$
@@
Assumption 2(i)-(ii) implies that $\sqrt{n}(\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1})\to_{p}\mathbf{0}$. Moreover, $\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|$, which is bounded in probability by Assumption 2(iv). It follows that $T_{1,n}\to_{p}\mathbf{0}$. For $T_{2,n}$, note that $\mathbb{E}[\mathbf{q}_{i}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}]=\mathbf{0}$ by independence of $\mathbf{q}_{i}$ and $\mathbf{x}_{i}$ conditional on $(C_{i},\bm{\theta}_{i})$. Let $\mathbf{X}_{i}=\mathbf{q}_{i}\left(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right)^{T}$ and $D$ the dimension of $\mathbf{q}_{i}$. Then
@@
First consider the denominator term. By Lemma 5 and condition (7), we have $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}]$. Hence, by Assumption 2(iv),
$\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}]^{-1}.$ (24)
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;(Y_{i}-\hat{\bm{\xi}}_{i}^{T}\bm{\psi})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}(\bm{\xi}_{i}-\hat{\bm{\xi}}_{i})^{T}\bm{\psi}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\varepsilon_{i}=:T_{1,n}+T_{2,n}.$
@@
$T_{1,n}\to_{p}\left[\begin{array}[]{c}-\kappa\,\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$ (25)
@@
$=\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|C_{i},\bm{\theta}_{i}\right]\frac{1}{C_{i}}\mathrm{tr}\left\{\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right\}\right]$
$=\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\varepsilon_{i}^{2}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right]\to 0.$
@@
Note that $\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ is finite by Assumption 2(iv)-(v). It follows by the central limit theorem that
$T_{2,n}\to_{d}N\left(\bm{0},\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]\right).$ (26)
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right].$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\left(\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;{\bm{\xi}}_{i}^{T}\right)\\ +\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2})\hat{\bm{\xi}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=:T_{3,n}+T_{4,n}+T_{5,n}.$
Evidently, $T_{3,n}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ by the LLN. Now consider $T_{4,n}$. By construction, we have
$T_{4,n}=\left[\begin{array}[]{cc}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T})&\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})\mathbf{q}_{i}^{T}\\ \frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}&\mathbf{0}\end{array}\right].$